# Mixture-of-Linear-Experts for Long-term Time Series Forecasting

### v2

@@ Abstract
Long-term time series forecasting (LTSF) aims to predict future values of a time series given the past values. The current state-of-the-art (SOTA) on this problem is attained in some cases by linear-centric models, which primarily feature a linear mapping layer. However, due to their inherent simplicity, they are not able to adapt their prediction rules to periodic changes in time series patterns. To address this challenge, we propose a Mixture-of-Experts-style augmentation for linear-centric models and propose Mixture-of-Linear-Experts (MoLE). Instead of training a single model, MoLE trains multiple linear-centric models (i.e., experts) and a router model that weighs and mixes their outputs. While the entire framework is trained end-to-end, each expert learns to specialize in a specific temporal pattern, and the router model learns to compose the experts adaptively. Experiments show that MoLE reduces forecasting error of linear-centric models, including DLinear, RLinear, and RMLP, in over 78% of the datasets and settings we evaluated. By using MoLE existing linear-centric models can achieve SOTA LTSF results in 68% of the experiments that PatchTST reports and we compare to, whereas existing single-head linear-centric models achieve SOTA results in only 25% of cases. Additionally, MoLE models achieve SOTA in all settings for the newly released Weather2K datasets.
@@ 1 Introduction
However, real-world time series are usually non-stationary. For example, traffic patterns change on different days of the week. Due to the inherent simplicity of linear-centric models, it is difficult for them to capture these patterns.
@@
  • We propose MoLE, a mixture-of-experts approach that can augment any linear-centric LTSF model to capture non-stationary temporal patterns.
  • We demonstrate through comprehensive empirical evaluations that MoLE consistently improves upon existing linear-centric methods, including those that are currently SOTA for the LTSF problem. Concretely, we observe when MoLE was applied to DLinear, it enhanced its performance in 32/44 experimental settings (73% of experiments). Similarly, with the RLinear and RMLP models, integrating MoLE resulted in improvements in 38/44 scenarios (86%) and 33/44 scenarios (75%) respectively. We find that among the 28 datasets and settings that PatchTST reports and we compare to, MoLE allows linear-centric models to achieve SOTA results in 19 cases (68%), whereas without MoLE (single-head), they are SOTA in 7 settings (25%). For the Weather2K datasets, MoLE-linear-centric models achieve SOTA in all settings.
@@ 2 Related Work
Gruver et al. (2023) recently introduced the LLMTime model, which uses large language models as a zero-shot time series predictor. We were unable to compare to this model due to the high volume and prohibitive cost of API calls to proprietary LLMs. Their quantitative results also did not include exact numbers, but were listed as bar charts, so we could not compare against reported results either.
@@ 4.2 Toy Dataset
To show the intuitive effect of this approach, we first apply it to a toy dataset. We constructed the dataset such that from Monday to Thursday, the values adhere to a sinusoid function with frequency $f$. However, from Friday to Sunday, the values follow a sinusoid function with a doubled frequency of $2f$. Both have Gaussian noise added at each time step. This dataset exhibits weekly periodicity. The exact distribution is provided in B.1 in Supplementary Materials, and the first two weeks of the datasets are visualized in Figure 3.
@@
For this experiment, we use RLinear as the backbone. The setup includes an input sequence length of 24 (1 day), prediction length of 24 (1 day), batch size of 128, and an initial learning rate of 0.005. We restrict ourselves to only 2 heads for the MoLE-RLinear model in this experiment.
@@ 5.1 Experimental Setup
Table 1 provides a brief overview of these datasets. For the ETT datasets, data is split into training, validation, and test sets in a temporal order with a ratio of 6:2:2. For all other datasets, the split ratio is 7:1:2. The split ratios we have employed in our experiments are consistent with the methodologies used in prior work.
Table 1: Overview of Datasets
Size Granu-
Dataset (Timestamps larity
$\times$×Channels)
ETTh1,h2 17420 (725.83 days) $\times$ 7 1 hour
ETTm1,m2 69680 (725.83 days) $\times$ 7 15 minutes
Weather 52696 (365.86 days) $\times$ 21 10 minutes
Electricity 26304 (1096 days) $\times$ 321 1 hour
Traffic 17544 (731 days) $\times$ 336 1 hour
Weather2K 40896 (1704 days) $\times$ 20 1 hour
@@
We applied MoLE to three recent linear-centric LTSF models: DLinear (Zeng et al., 2023), RLinear (Li et al., 2023), and RMLP (Li et al., 2023). We refer to the original models as original and the models enhanced by our method as MoLE. Additionally, we evaluated the state-of-the-art transformer model, PatchTST (Nie et al., 2022). We assessed all models across every dataset, reproducing prior results of linear-centric models ourselves. Since PatchTST involved extensive hyperparameter search, we decided to use the author-reported results. Since PatchTST was not evaluated in the original paper on the newly released Weather2K dataset, we do not report PatchTST’s results for Weather2K. In our comparisons, we used the results of supervised PatchTST/64 and PatchTST/42. PatchTST/64 means the number of input patches is 64, which uses the look-back window $L=512$. PatchTST/42 means the number of input patches is 42, which has the default look-back window $L=336$. Our results on linear-centric models were mostly similar to reported results, so we use our computed values for our evaluation. In Appendix C (Supplementary Material), we give a complete table that compares the reported results from prior work to the values we obtained experimentally, and provide additional discussion of these variations. Additionally, we include the outcomes of further hyperparameter tuning to assess their impact on model performance in Appendix C.3.
@@
As in existing studies, we adopt the widely-used evaluation metric, mean squared error (MSE), which has been utilized in the works of Wu et al. (2021), Zhou et al. (2021), Zhang and Yan (2022), Zeng et al. (2023), and Li et al. (2023). The metric is defined as:
$\operatorname{MSE}=\frac{1}{N}\sum_{i=1}^{N}(\mathbf{Y}_{i}-\hat{\mathbf{Y}}_{i})^{2}$
where $N$ represents the total number of samples, $\mathbf{Y}_{i}$ indicates the i-th prediction, and $\hat{\mathbf{Y}}_{i}$ is the corresponding observation (ground truth).
@@
Head dropout. To combat overfitting, we incorporated head dropout into MoLE. Throughout training, the weights originating from the mixing layer are randomly dropped at a rate of $r$. The retained weights are adjusted to ensure their combined sum is 1. We tested two rates: 0 (head dropout deactivated) and 0.2. We also provide a detailed analysis and comprehensive results on how the choice of dropout rate impacts multi-head model performance in Appendix D.2.
@@ 5.2 Comparison with Single-Head Models
Table 3: Comparison between original (single-head) and enhanced (multi-head MoLE) DLinear, RLinear, and RMLP models. Prediction length $\in\{96,192,336,720\}$. The values reported are MSE loss. A lower value indicates a better prediction. The better results among a linear-centric method and its MoLE variant is highlighted in bold. The cells shown in blue indicate the current SOTA for a given dataset and prediction length. For all three linear-centric models, MoLE improves forecasting error in over 78% of the datasets and settings we evaluated. Blank entries indicate datasets and prediction lengths that were not reported by the authors of PatchTST.
@@ 5.3 Ablations and Design Choices
Surprisingly, RandomIn and RandomOut occasionally produce losses lower than TimeIn, which uses timestamp embeddings. This is consistent with the of findings of Roller et al. (2021) on MoE with hash layers. We explore why in the next subsection. For now, we comment on the robustness of TimeIn. As shown in Table 4, across all experiments, TimeIn had the lowest loss in 25 instances, surpassing RandomIn and RandomOut, which achieved this 11 and 8 times, respectively.
@@
We next try to understand when and why random inputs (and even outputs!) can help linear-centric models. We propose two possible effects contributing to this observation: Effect 1: RandomIn and RandomOut may have a regularization effect, akin to dropout (Baldi and Sadowski, 2013), which prevents the network from overfitting head weights to the training data. This has the effect of making all heads better. Effect 2: Suppose time series can be roughly divided into regimes with different temporal patterns (e.g., weekdays vs weekends). For long input sequences, the sequence may traverse multiple regimes. Hence, no single head will be able to capture the full dynamics, regardless of start time. However, as the input sequence lengthens, the linear layer has more past data to consider, and some patterns in this data can hint at upcoming changes in the series. Hence, we predict that the effect of conditioning on start time should be less beneficial. Conversely, as the input sequence length is reduced, the effect of input time should be more beneficial.
@@
We finally present the results of our experiments to determine the batch size for our prior experiments. Figure 6 plots how the final test loss varies with batch size, for four datasets. For these experiments, we used DLinear as the base model, with an input sequence length of 336, prediction length of 96, five heads, and an initial learning rate of 0.05. We observe that in general, test losses are lower when batch sizes are either very small or (in some cases) large (relative to intermediate sizes ranging from 16 to 128). This mid-range of batch sizes exhibits poor generalization, with low training loss but high test loss, similar to observations in Keskar et al. (2016). For robustness, we used a batch size of 8. Comprehensive results and a more detailed analysis of batch size effects are available in the Appendix D.1.
@@ 6 Conclusion and Future Work
We propose MoLE, a technique based on mixture-of-experts that can be applied out-of-the-box to existing linear-centric models to improve performance in the LTSF problem. Our results strongly suggest that MoLE gives significant and robust gains across a wide range of datasets and prediction settings. An important open question is precisely characterizing high-dimensional datasets to understand exactly when MoLE will help. Although we hypothesize that this question relates to the degree and patterns of nonstationarity in the time series dataset, it is currently unclear how to predict how many heads are needed.
@@ B.2 More detailed results from the toy dataset experiments
Figure 6(a) shows how the mixing weights change over a continuous 2-week period. The mixing layer captures the temporal patterns in the data, applying different heads for different frequencies. Figure 6(b) is a heatmap representing the weight distribution of the linear layer in the single-head model. Using one head to predict the time series, the model derives a compromised weight set that underperforms, as shown in Figure 4. However, when employing two heads, the resulting models—whose linear layer weights are depicted in Figures 6(c) and 6(d)—efficiently learn the two frequency patterns and can yield more accurate predictions.
Appendix C Additional Main Results
C.1 Comparing Original and Our Replicated Experiments.
In this section, we compare previously-reported results with our reproduction of prior linear-centric results (using their code).
@@ Appendix C Additional Main Results
C.2 Results with Different Random Seeds
@@
C.3 Results with Additional Hyperparameter Tuning
@@
Table 8: Comparison between original (single-head) and enhanced (multi-head MoLE) DLinear, RLinear, and RMLP models with added hyperparameter values. Prediction length $\in\{96,192,336,720\}$. The values reported are MSE loss. A lower value indicates a better prediction. The better results among a linear-centric method and its MoLE variant is highlighted in bold.
@@
Appendix D Additional Ablations
Based on the results of the full hyperparameter tuning shown in Appendix C.3, we conducted two additional ablations to further discuss how the choices of batch sizes and head dropout rates impact model performance. The plots that follow show curves which are the average test losses from three different seeds. These test losses are chosen from the hyperparameter configurations that yielded the lowest validation losses, while keeping the parameter on the x-axis constant. The shaded regions around each curve indicate the range of $\pm$ one standard deviation.
D.1 Impact of Batch Sizes on Model Performance
Figure 8 demonstrates the impact of varying batch sizes on the performance of both single-head and MoLE DLinear, RLinear, and RMLP models. For smaller datasets, such as the ETT series, there is a trend similar to an inverse U-shape, suggesting that very small or very large batch sizes may enhance model performance. In contrast, for larger datasets, particularly electricity and traffic, larger batch sizes appear to benefit single-head models, while MoLE models show optimal performance with mid-range to small batch sizes. Importantly, in large datasets, MoLE models consistently outperform single-head models, even when the latter are given the advantage of larger batch sizes. This trend is also observable in Table 8, where comprehensive hyperparameter tuning does not enable single-head models to outperform MoLE in large datasets. These results confirm that the batch size of 8, used in our main experiments, is sufficient to effectively represent each model’s capabilities.

### v3

@@ Abstract
Long-term time series forecasting (LTSF) aims to predict future values of a time series given the past values. The current state-of-the-art (SOTA) on this problem is attained in some cases by linear-centric models, which primarily feature a linear mapping layer. However, due to their inherent simplicity, they are not able to adapt their prediction rules to periodic changes in time series patterns. To address this challenge, we propose a Mixture-of-Experts-style augmentation for linear-centric models and propose Mixture-of-Linear-Experts (MoLE). Instead of training a single model, MoLE trains multiple linear-centric models (i.e., experts) and a router model that weighs and mixes their outputs. While the entire framework is trained end-to-end, each expert learns to specialize in a specific temporal pattern, and the router model learns to compose the experts adaptively. Experiments show that MoLE reduces forecasting error of linear-centric models, including DLinear, RLinear, and RMLP, in over 78% of the datasets and settings we evaluated. By using MoLE existing linear-centric models can achieve SOTA LTSF results in 68% of the experiments that PatchTST reports and we compare to, whereas existing single-head linear-centric models achieve SOTA results in only 25% of cases.
@@ 1 INTRODUCTION
However, real-world time series often exhibit seasonal variations and non-stationarity. For example, traffic patterns change on different days of the week. Due to the inherent simplicity of linear-centric models, it is difficult for them to capture these patterns.
@@
  • We propose MoLE, a mixture-of-experts approach that can augment any linear-centric LTSF model to capture non-stationary, seasonal temporal patterns.
  • We demonstrate through comprehensive empirical evaluations that MoLE improves upon existing linear-centric methods, including those that are currently SOTA for the LTSF problem. Concretely, we observe when MoLE was applied to DLinear, it enhanced its performance in 32/44 experimental settings (73% of experiments). Similarly, with the RLinear and RMLP models, integrating MoLE resulted in improvements in 38/44 scenarios (86%) and 33/44 scenarios (75%) respectively. We find that among the 28 datasets and settings that PatchTST reports and we compare to, MoLE allows linear-centric models to achieve SOTA results in 19 cases (68%), whereas without MoLE (single-head), they are SOTA in 7 settings (25%).
@@ 2 RELATED WORK
Gruver et al. (2023) recently introduced the LLMTime model, which employs large language models as a zero-shot predictor for time series. We were unable to compare with this model due to the prohibitive costs of API calls to proprietary LLMs. Additionally, their quantitative results were presented as bar charts without exact numbers, preventing a direct comparison of reported outcomes.
@@ 4.2 Toy Dataset
To show the intuitive effect of this approach, we first apply it to a toy dataset. We constructed the dataset such that from Monday to Thursday, the values adhere to a sinusoid function with frequency $f$. From Friday to Sunday, the values follow a sinusoid function with a doubled frequency of $2f$. Both have Gaussian noise added at each time step. This dataset exhibits weekly periodicity. The exact distribution is provided in B.1 in Supplementary Materials, and the first two weeks of the datasets are visualized in Figure 3.
@@
For this experiment, we use RLinear as the backbone. The setup includes an input sequence length of 24 (1 day), a prediction length of 24 (1 day), a batch size of 128, and an initial learning rate of 0.005. We restrict ourselves to only 2 heads for the MoLE-RLinear model in this experiment.
@@ 5.1 Experimental Setup
Table 4 in Supplementary Materials provides a brief overview of these datasets. For the ETT datasets, data is split into training, validation, and test sets in a temporal order with a ratio of 6:2:2. For all other datasets, the split ratio is 7:1:2. The split ratios we have employed in our experiments are consistent with the methodologies used in prior work.
@@
We applied MoLE to three recent linear-centric LTSF models: DLinear (Zeng et al., 2023), RLinear (Li et al., 2023), and RMLP (Li et al., 2023). We refer to the original models as original and the models enhanced by our method as MoLE. Additionally, we evaluated the state-of-the-art transformer model, PatchTST (Nie et al., 2022). We evaluated all models on each dataset and reproduced results for linear-centric models. For PatchTST, we relied on author-reported outcomes. PatchTST’s results for the newly released Weather2K dataset are not included, as it was not assessed in the original study. In our comparisons, we used the results of supervised PatchTST/64 and PatchTST/42. PatchTST/64 means the number of input patches is 64, which uses the look-back window $L=512$. PatchTST/42 means the number of input patches is 42, which has the default look-back window $L=336$. Our results on linear-centric models were mostly similar to reported results, so we use our computed values for our evaluation. In Appendix D (Supplementary Material), we give a complete table that compares the reported results from prior work to the values we obtained experimentally, and provide additional discussion of these variations. Additionally, we include the outcomes of further hyperparameter tuning to assess their impact on model performance in Appendix D.3.
@@
As in existing studies, we adopt the widely-used evaluation metric, mean squared error (MSE), which has been utilized in the works of Wu et al. (2021), Zhou et al. (2021), Zhang and Yan (2022), Zeng et al. (2023), and Li et al. (2023). The metric is defined as: $\operatorname{MSE}=\frac{1}{N}\sum_{i=1}^{N}(\mathbf{Y}_{i}-\hat{\mathbf{Y}}_{i})^{2}$ where $N$ represents the total number of samples, $\mathbf{Y}_{i}$ indicates the i-th prediction, and $\hat{\mathbf{Y}}_{i}$ is the corresponding observation (ground truth).
@@
Head dropout. To combat overfitting, we incorporated head dropout into MoLE. Throughout training, the weights originating from the mixing layer are randomly dropped at a rate of $r$. The retained weights are adjusted to ensure their combined sum is 1. We tested two rates: 0 (head dropout deactivated) and 0.2. We also provide a detailed analysis and comprehensive results on how the choice of dropout rate impacts multi-head model performance in Appendix E.2.
@@ 5.2 Comparison with Single-Head Models
Table 2: Comparison between original (single-head) and enhanced (multi-head MoLE) DLinear, RLinear, and RMLP models. Prediction length $\in\{96,192,336,720\}$. The values reported are MSE loss. A lower value indicates a better prediction. The better results among a linear-centric method and its MoLE variant are highlighted in bold. The cells shown in blue indicate the current SOTA for a given dataset and prediction length. For all three linear-centric models, MoLE improves forecasting error in over 78% of the datasets and settings we evaluated. Entries marked with ‘–’ indicate datasets and prediction lengths not reported in previous work.
@@ 5.3 Ablations and Design Choices
Surprisingly, RandomIn and RandomOut occasionally produce losses lower than TimeIn, which uses timestamp embeddings. This is consistent with the findings of Roller et al. (2021) on MoE with hash layers. We explore why in the next subsection. For now, we comment on the robustness of TimeIn. As shown in Table 3, across all experiments, TimeIn had the lowest loss in 25 instances, surpassing RandomIn and RandomOut, which achieved this 11 and 8 times, respectively.
@@
We next try to understand when and why random inputs and outputs can help linear-centric models. We propose two possible effects contributing to this observation: Effect 1: RandomIn and RandomOut may have a regularization effect, akin to dropout (Baldi and Sadowski, 2013), which prevents the network from overfitting head weights to the training data. This has the effect of making all heads better. Effect 2: Suppose time series can be roughly divided into regimes with different temporal patterns (e.g., weekdays vs weekends). For long input sequences, the sequence may traverse multiple regimes. Hence, no single head will be able to capture the full dynamics, regardless of start time. However, as the input sequence lengthens, the linear layer has more past data to consider, and some patterns in this data can hint at upcoming changes in the series. Hence, we predict that the effect of conditioning on start time should be less beneficial. Conversely, as the input sequence length is reduced, the effect of input time should be more beneficial.
@@
We finally present the results of our experiments to determine the batch size for our prior experiments. Figure 6 plots how the final test loss varies with batch size, for four datasets. For these experiments, we used DLinear as the base model, with an input sequence length of 336, prediction length of 96, five heads, and an initial learning rate of 0.05. We observe that in general, test losses are lower when batch sizes are either very small or (in some cases) large (relative to intermediate sizes ranging from 16 to 128). This mid-range of batch sizes exhibits poor generalization, with low training loss but high test loss, similar to observations in Keskar et al. (2016). For robustness, we used a batch size of 8. Comprehensive results and a more detailed analysis of batch size effects are available in Appendix E.1.
@@ 6 CONCLUSION AND FUTURE WORK
We propose MoLE, a technique based on mixture-of-experts that can be applied out-of-the-box to existing linear-centric models to improve performance in the LTSF problem. Our results strongly suggest that MoLE gives significant and robust gains across a wide range of datasets and prediction settings. An important open question is precisely characterizing high-dimensional datasets to understand exactly when MoLE will help. Although we hypothesize that this question relates to the variability and complexity of patterns in the time series dataset, it is currently unclear how to predict how many heads are needed.
@@ Acknowledgments
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. Yes
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. Yes
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Yes
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. Not Applicable
    2. (b)
      Complete proofs of all theoretical results. Not Applicable
    3. (c)
      Clear explanations of any assumptions. Not Applicable
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). Yes
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). Yes
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). Yes
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). Yes
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. Yes
    2. (b)
      The license information of the assets, if applicable. Not Applicable
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. Not Applicable
    4. (d)
      Information about consent from data providers/curators. Not Applicable
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not Applicable
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. Not Applicable
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not Applicable
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not Applicable
Mixture-of-Linear-Experts for Long-term Time Series Forecasting:
@@ B.2 More detailed results from the toy dataset experiments
Figure 7(a) shows how the mixing weights change over a continuous 2-week period. The mixing layer captures the temporal patterns in the data, applying different heads for different frequencies. Figure 7(b) is a heatmap representing the weight distribution of the linear layer in the single-head model. Using one head to predict the time series, the model derives a compromised weight set that underperforms, as shown in Figure 4. However, when employing two heads, the resulting models—whose linear layer weights are depicted in Figures 7(c) and 7(d)—efficiently learn the two frequency patterns and can yield more accurate predictions.
@@
Appendix C Overview of Real-World Datasets
Table 4 provides a brief overview of the real-world datasets evaluated.
Table 4: Overview of Datasets
Size Granu-
Dataset (Timestamps larity
$\times$×Channels)
ETTh1,h2 17420 (725.83 days) $\times$ 7 1 hour
ETTm1,m2 69680 (725.83 days) $\times$ 7 15 minutes
Weather 52696 (365.86 days) $\times$ 21 10 minutes
Electricity 26304 (1096 days) $\times$ 321 1 hour
Traffic 17544 (731 days) $\times$ 336 1 hour
Weather2K 40896 (1704 days) $\times$ 20 1 hour
Appendix D Additional Main Results
D.1 Comparing Original and Our Replicated Experiments.
In this section, we compare previously reported results with our reproduction of prior linear-centric results (using their code).
@@ Appendix D Additional Main Results
D.2 Results with Different Random Seeds
@@
D.3 Results with Additional Hyperparameter Tuning
@@
Table 8: Comparison between original (single-head) and enhanced (multi-head MoLE) DLinear, RLinear, and RMLP models with added hyperparameter values. Prediction length $\in\{96,192,336,720\}$. The values reported are MSE loss. A lower value indicates a better prediction. The better results among a linear-centric method and its MoLE variant are highlighted in bold.
@@
Appendix E Additional Ablations
Based on the results of the full hyperparameter tuning shown in Appendix D.3, we conducted two additional ablations to further discuss how the choices of batch sizes and head dropout rates impact model performance. The plots that follow show curves which are the average test losses from three different seeds. These test losses are chosen from the hyperparameter configurations that yielded the lowest validation losses, while keeping the parameter on the x-axis constant. The shaded regions around each curve indicate the range of $\pm$ one standard deviation.
E.1 Impact of Batch Sizes on Model Performance
Figures 8, 9 and 10 demonstrate the impact of varying batch sizes on the performance of both single-head and MoLE DLinear, RLinear, and RMLP models. For smaller datasets, such as the ETT series, there is a trend similar to an inverse U-shape, suggesting that very small or very large batch sizes may enhance model performance. In contrast, for larger datasets, particularly electricity and traffic, larger batch sizes appear to benefit single-head models, while MoLE models show optimal performance with mid-range to small batch sizes. Importantly, in large datasets, MoLE models consistently outperform single-head models, even when the latter are given the advantage of larger batch sizes. This trend is also observable in Table 8, where comprehensive hyperparameter tuning does not enable single-head models to outperform MoLE in large datasets. These results confirm that the batch size of 8, used in our main experiments, is sufficient to effectively represent each model’s capabilities.
E.2 Impact of Head Dropout Rates on MoLE Performance